# Learning Dynamic Representations via An Optimally-Weighted Maximum Mean Discrepancy Optimization Framework for Continual Learning

### v4

@@ 4.4 Analysis Results
With the smaller buffer size (200), OWMMD achieves the highest accuracy in both scenarios (67.74% Class-IL, 92.97% Task-IL), outperforming all compared methods. Notably, its Class-IL accuracy surpasses the second-best DER$+$Refresh by 2.35%, while maintaining a 0.17% Task-IL advantage. This demonstrates superior capability in memory-constrained environments where methods like GEM (25.54% Class-IL) and A-GEM (20.04% Class-IL) show significant degradation.
@@ 4.5 Ablation Study
Table 8: Ablation study on CIFAR-10 with buffer size 500. "First $n$" indicates that only the first $n$ layers of the backbone are utilized for matching and MMD calculation, while "Last $n$" refers to using only the last $n$ layers. "ALL" indicates that all backbone layers are used. AROT denotes Adaptive Regularization Optimization Term.

### v5

@@ 5.5 Analysis Results
With the smaller buffer size (200), OWMMD achieves the highest accuracy in both scenarios (67.74% Class-IL, 92.97% Task-IL), outperforming all compared methods. Notably, its Class-IL accuracy surpasses the second-best DER+++Refresh by 2.35%, while maintaining a 0.17% Task-IL advantage. This demonstrates superior capability in memory-constrained environments where methods like GEM (25.54% Class-IL) and A-GEM (20.04% Class-IL) show significant degradation.
@@ 5.6 Ablation Study
Table 9: Ablation study on CIFAR-10 with buffer size 500. "First nn" indicates that only the first $n$ layers of the backbone are utilized for matching and MMD calculation, while "Last nn" refers to using only the last $n$ layers. "ALL" indicates that all backbone layers are used. AROT denotes Adaptive Regularization Optimization Term.
@@ 6 Conclusion and Future Works
Task-Free Continual Learning (TFCL) presents a more complex paradigm, characterized by the absence of explicit task delineation and boundaries throughout both training and evaluation stages. Given that the proposed OWMMD framework utilizes reservoir sampling for dynamic memory buffer updates, its memory management strategy is inherently compatible with TFCL settings. Nevertheless, a notable constraint in adapting OWMMD to TFCL lies in the reliance of the MLFMM component on task-specific information to archive the current model as a teacher module for regularization loss computation. To overcome this limitation, we would like to develop a straightforward yet effective task-boundary detection technique that leverages the emergence of novel categories as indicators of task transitions. This direction will be systematically explored in our future research. We have added the discussion in the conclusion and limitation section of the revised version.